Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by danjuma » Sat Dec 04, 2010 10:13 pm
A basket contains 5 apples of which 1 is spoiled and the rest are good. If Henry is to select 2 apples from the basket simultaneously and at random, what is the probability that 2 apples selected will include the spoiled apple?

a.1/5 b.3/10 c.2/5 d.1/2 e.3/5
Join the discussion
Source: — Problem Solving |

by beat_gmat_09 » Sat Dec 04, 2010 10:30 pm
Two apples are drawn simultaneously, so order does not matter (both are picked at one time).
There is 1 spoiled apple, 1 good apple, total = 2
Possible outcomes = 5
Probability = 2/5
Hope is the dream of a man awake
Join the discussion

by goyalsau » Sun Dec 05, 2010 1:52 am
beat_gmat_09 wrote:Two apples are drawn simultaneously, so order does not matter (both are picked at one time).
There is 1 spoiled apple, 1 good apple, total = 2
Possible outcomes = 5
Probability = 2/5
it should be 2/5 or 40%
Saurabh Goyal
[email protected]
-------------------------


EveryBody Wants to Win But Nobody wants to prepare for Win.
Join the discussion

by Brian@VeritasPrep » Tue Dec 07, 2010 12:34 pm
Hey guys,

Just one point that may be worth a chime-in on these "simultaneous" probability problems. I think a lot of times the word "simultaneous" throws people off, but when you think about it there's no such thing as perfectly "simultaneous" if you take it all the way down to a nanosecond level, right? At some point your finger touches one apple before the other. So you can solve something like this methodically by using both sequences:

Spoiled, then Not Spoiled (1/5 * 4/4)

Not Spoiled, then Spoiled (4/5 * 1/4)

And then adding them together (4/20 + 4/20 = 8/20 = 2/5)

Or you can think through logically: If he picks 2 and each has a 1/5 chance of being "the spoiled one", then he has two 1/5 chances of a spoiled apple, or a 2/5 total chance.


Just thought it was worth pointing out...the word "simultaneous" in these problems carries a lot more intimidation than it's really worth...
Brian Galvin
GMAT Instructor
Chief Academic Officer
Veritas Prep

Looking for GMAT practice questions? Try out the Veritas Prep Question Bank. Learn More.
Join the discussion